OpenStudy (anonymous):

Solve the initial value problem explicitly: dx/dt = (8/t) - (1/t^4) + 8 and x = 6 when t = 1

OpenStudy (sweetburger):

what are you solving for and is this the entire question?

OpenStudy (anonymous):

Yes i have to find the integral of that equation and find C

OpenStudy (anonymous):

Any help?

OpenStudy (anonymous):

\[dx=\left[ \frac{ 8 }{ t }-\frac{ 1 }{ t^4 }+8 \right]dt\] integrate both sides \[x=8 \int\limits \frac{ 1 }{ t }dt- \int\limits t ^{-4}dt+8\int\limits dt+c\] \[x=8 \ln \left| t \right|-\frac{ t ^{-3} }{ -3 }+8t+c\] find c when x=6,t=1 then substitute the value of c

OpenStudy (anonymous):

How do you know t^-4 integrates to t^-3/-3?

OpenStudy (anonymous):

I got C = -7/3

OpenStudy (anonymous):

\[\int\limits x^n dx =\frac{ x ^{n+1} }{ n+1 }\]

OpenStudy (anonymous):

\[\int\limits t^{-4}dt=\frac{ t ^{-4+1} }{ -4+1 }=\frac{ t ^{-3} }{ -3 }\]

OpenStudy (anonymous):

So would the integral of -(24/x^7) = x^-6/-144?

OpenStudy (anonymous):

\[\int\limits \frac{- 24 }{ x^7 }dx=-24*\frac{ x ^{-6} }{ -6 }\] \[=\frac{ 4 }{ x^6 }\]

OpenStudy (anonymous):

Why does the variable move to the bottom?